40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 6320     Entry time: Sat Feb 25 00:37:42 2012
Author: kiwamu 
Type: Update 
Category: SUS 
Subject: oplev spectra during PRMI lock 

Somehow the angular stability of the central part have not been so great.

Also the angular motions look fluctuating a lot and they seem to be related with the glitches.

I took the oplev spectra when the PRMI is locked and unlocked to see whether if something obviously crazy is going on or not.

They seem ok to me except that the PRM pitch shows an extra bump at around 2-3 Hz when the PRMI is locked. But I don't think it's prominent.

 


- The attached files show the oplev spectra. When the PRMI is locked the PRM and both ITMs are under the length control.

(red) pitch when PRMI is locked

(blue) yaw when PRMI is locked

(orange) pitch without any length controls

(cyan) pitch without any length controls

 

Attachment 1: oplev_PRMI.pdf  434 kB  | Hide | Hide all
oplev_PRMI.pdf oplev_PRMI.pdf oplev_PRMI.pdf oplev_PRMI.pdf
ELOG V3.1.3-